OpenStudy (anonymous):

How am I able to simplify: (7/2x) + (3/5x), step by step, please!

OpenStudy (anonymous):

(7/2+3/5)x = (7*5 +3*2)x/10 = 41x/10

OpenStudy (anonymous):

the top guy is correct idk if there is any oother way to simplify the final answer, i hope ur teacher would accept that as a final answer

OpenStudy (anonymous):

Yes, thank you guys!

OpenStudy (anonymous):